OpenStudy (anonymous):

the sum of two numbers is 40. the sum of the smaller and 4 times the larger is 115. find the numbers

OpenStudy (anonymous):

x+y=40 x+4y=115

myininaya (myininaya):

x=40-y x=115-4y so 40-y=115-4y 40=115-3y 40-115=-3y -3y=-75 y=25 so x=40-25=15